OpenStudy (anonymous):

In the below system, solve for y in the first equation. 3x + y = 9 4x – 3y = 6

OpenStudy (anonymous):

y = 9 - 3x

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

subtract 3x from both sides

OpenStudy (anonymous):

Can you draw it so I can see what your talking about?

OpenStudy (anonymous):

Actually, if this is a system of equations, bob's answer isn't completely correct. What you want to do is solve for x in the first equation: 3x + y = 9 ---> 3x = 9 - y ---> x = (9-y)/3 now plug this value into your second equation: --> 4[(9-y)/3] - 3y = 6 (plugged x = (9-y)/3 in for x) and solve for y: 4*(9/3 - y/3) - 3y = 6 36/3 - 4y/3 - 3y = 6 (distributed the 4) 4y/3 - 3y = 6 - 12 (simplified 36/3 and took it to the other side) 4y - 9y = 18 - 36 (multiplied everything by 3 to make things simpler) -5y = -18 (divide by -5) y = -18/-5 y = 18/5 or 3.6 so remember: solve for one of the variables in the first equation then plug it into the second equation to get your answer

OpenStudy (anonymous):

Thank you!!! <3

OpenStudy (anonymous):

ur welcome :)